OpenStudy (anonymous):

Let P be a point on the curve x=t^2 , y= 1/t. If the tangent to the curve at P meets the x- and y- axes at A and B respectively, prove that PA=2BP

OpenStudy (anonymous):

First, I differentiate x and y x'=2t y'=1/t^2

OpenStudy (anonymous):

Then, I get 1/t^2 / 2t = 1/2t^3 Now what do I do?

OpenStudy (anonymous):

\[={1\over 2t^3}\]

OpenStudy (dumbcow):

y' = -1/t^2

OpenStudy (anonymous):

OH, Yes, I forgot.. so, -1/2t^3

OpenStudy (anonymous):

How do I go on from here?

OpenStudy (dumbcow):

well you know the slope of tangent line at P now you need equation of line to find intercepts A and B i think we need to get it in terms of x or unparamaterize it let me play with it a bit

OpenStudy (anonymous):

Hm... but there's no value for t, is there?

OpenStudy (anonymous):

wouldn't the equation be y= 1/2t^3 +1/2t?

OpenStudy (anonymous):

\[y= (1/2t^3)x +1/2t\]

OpenStudy (dumbcow):

ok i think i got it, i was wrong you can leave it terms of t you are on the right track ^^ but the y_intercept is wrong point P is (t^2, 1/t) dy/dx = m = -1/2t^3 tangent line \[y -\frac{1}{t} =- \frac{1}{2t^{3}}(x-t^{2})\] \[y =- \frac{1}{2t^{3}}x +\frac{3}{2t}\] plug in 0 for x and y to get intercepts A and B \[A = (3t^{2},0)\] \[B = (0,\frac{3}{2t})\] to finish proof use distance formula

OpenStudy (anonymous):

Ah Ok. Yes, I keep forgetting to add the - sign.

OpenStudy (dumbcow):

@ishaan94 is there another way to do the problem

OpenStudy (anonymous):

That's the way it says to do it in my textbook, so I think I should stick with that method :)

OpenStudy (dumbcow):

oh ok

OpenStudy (anonymous):

I don't think so, the only method I know is the one you did.